OpenStudy (anonymous):

Fan and Medal -2x + 8 = -16 x = ? Please Show Steps

OpenStudy (shinalcantara):

transpose 8 to the right side of the equation. start from there

OpenStudy (anonymous):

I don't understand.

OpenStudy (ahsome):

\[-2x+8=-16\]\[-2x=-16-8\]\[-2x=-24\]\[x=\frac{-24}{-2}\]\[x=12\]

OpenStudy (shinalcantara):

Ahsome answered it for you.. i hope you understand how it's done.

OpenStudy (ahsome):

If you don't understand, please ask :) We won't bite

OpenStudy (anonymous):

yea thats what i meant when i said show your steps thx anyways

OpenStudy (ahsome):

No problem :D

OpenStudy (anonymous):

So basically you ended up adding 16 and 8 because they were both negatives right ?

OpenStudy (ahsome):

Well, we had \(-16\) and \(-8\) on the right hand side, right? So what is \(-16-8\)? \(-24\)

OpenStudy (ahsome):

Be Right Back. If you still have questions, I am pretty sure @shinalcantara can answer them for you till I get back

OpenStudy (anonymous):

you subtracted -16 from -8 i was thinking of multiplication i guess

OpenStudy (ahsome):

Oh, OK